OpenStudy (anonymous):

Use the shell method to find the volume of the solid generated by revolving the region y= sq. rt x, y=0, and x=4 about the line x=4.

OpenStudy (anonymous):

you need to begin by drawing or graping

OpenStudy (anonymous):



OpenStudy (anonymous):

That's a positive root x

OpenStudy (anonymous):

okay, now the inner radius is y=g(x) and the outer radius is given by y=f(x), in this case and the volume of the solid is given by v= intergral of pi[(f(x))^2-(g(x))^2]dx

OpenStudy (anonymous):

That would work if the ? was asking for washers. My professor is particular about following the directions. : | I've tried setting it up through the shell method and I keep on getting a negative answe.

OpenStudy (anonymous):

p(x)= (4-x) , h(x)= (x^1/2) is what I've done so far.

OpenStudy (anonymous):

I think te h(x) needs to be x^1/2 -4

OpenStudy (anonymous):

well, we need to look at a disk first

OpenStudy (anonymous):

lets say we has a disk, with with dx right

OpenStudy (anonymous):

now the surface are of this disk is (pi)r^2 which is equal to (pi)(sqrt x)^2

OpenStudy (anonymous):

oh, but you want shell method right

OpenStudy (anonymous):

disk method is so much easier

OpenStudy (anonymous):

I've done the disk method already. She wants both methods sadly.